# InterActHuman: Multi-Concept Human Animation with Layout-Aligned Audio Conditions

### v1

@@ 3.4 Multi-Concept Human-Centric Video Data Curation
We curate a large-scale, human-centric video dataset from a large internal video dataset, filtering out videos that are too short (<4 seconds) or contain too many salient humans (>3) as detected by a human pose detector dwpose . Unlike previous multi-concept video generation methods chen2025multi ; huang2025conceptmaster that rely on conventional pipelines like Grounding-DINO liu2023grounding and SAM kirillov2023segment for obtaining a single reference image per identity, our multi-stage pipeline leverages advanced vision-language models. First, each raw video is densely captioned using Qwen2-VL wang2024qwen2 (distilled from Gemini-2.0-Pro), generating fine-grained descriptions of the environment, subjects’ appearance, actions, expressions, interacting salient objects, and inter-subject interactions. Next, these descriptions are parsed by a zero-shot Gemini-2.0-Flash API to extract structured appearance phrases. For spatial supervision, we use Grounding-SAM2 ren2024grounded with the query person to produce accurate, temporally consistent masks, which are used both for extracting foreground reference images (with a white background) and as ground-truth for our mask predictor. Our corpus comprises over $2.6$M triplets of videos, per-frame masks, and captions, forming the foundation of InterActHuman.
@@ 4 Experiments
Baselines. Since ConceptMaster huang2025conceptmaster and Video-Alchemist chen2025multi are not open-sourced, we do not have access to their models nor test sets. Therefore, we primarily compare our method with recent multi-concept video customization models through their publicly available APIs or public models, evaluating performance from the perspectives of visual appearance and adherence to text prompts, including Vidu2.0 bao2024vidu , Pika2.1 pikaart2024 , Kling 1.6 kuaishou2024 and Phantom liu2025phantom , and audio-driven human video generation methods, including DiffTED diffted , DiffGest zhu2023taming + Mimiction zhang2024mimicmotion , CyberHost lin2024cyberhost , OmniHuman lin2025omnihuman and Kling 1.6 kuaishou2024 with lip synchronization.
@@
DiffTED diffted 0.926 - 0.769 - - - -
DiffGest. zhu2023taming +Mimic. zhang2024mimicmotion 0.496 23.409 0.833 - - - -
CyberHost lin2024cyberhost 6.627 24.733 0.884 8.974 4.011 2.856 54.797
Kling1.6 kuaishou2024 + Lip-sync. 4.449 46.490 0.826 8.401 4.716 3.444 33.555
OmniHuman lin2025omnihuman w/o mask 7.443 47.561 0.898 9.482 4.768 3.466 33.895
OmniHuman lin2025omnihuman w/ fixed mask - - - 7.068 4.690 3.369 40.239
@@
Metric Kling kuaishou2024 OmniHuman lin2025omnihuman Ours Pika pikaart2024 Phantom liu2025phantom Kling kuaishou2024 Vidu bao2024vidu Ours
@@
Vidu2.0 bao2024vidu 0.696 0.458 0.568 0.562 0.597 18.61 3.350 4.689
Pika2.1 pikaart2024 0.688 0.459 0.579 0.566 0.607 19.39 3.534 4.791
Kling1.6 kuaishou2024 0.659 0.420 0.552 0.547 0.582 18.38 3.487 4.787
Phantom liu2025phantom 0.703 0.476 0.589 0.573 0.615 17.73 3.404 4.812
@@
Evaluation Metrics. As audio and appearance are key characteristics of humans, we follow current state-of-the-art method lin2025omnihuman ; huang2025conceptmaster to evaluate audio-driven, multi-concept human-centric video generation. To comprehensively evaluate methods for this task, we consider five dimensions: 1) Concept fidelity: We leverage CLIP-I radford2021learning and DINO-I oquab2023dinov2 on subjects in the generated videos to assess whether they are aligned with the provided reference images. To get cropped subject images in output videos, we randomly sample 5 frames in each video and use Florence-2 xiao2024florence to detect the subject bounding boxes. We then crop the detected regions and compute the CLIP-I and DINO-I scores. We also employ Face-Arc deng2019arcface , Face-Cur huang2020curricularface and Face-Glink deng2019arcface to evaluate the fidelity of facial features if the concept is a human. 2) Prompt following: We utilize video-level CLIP wang2022internvideo to measure the similarity between the input text prompts and the resulting video content. 3) Visual quality: We utilize q-align wu2023q , a vision-language model, for no-reference image quality assessment (IQA) and aesthetics score estimation (ASE). 4) Audio synchronization and human pose diversity: For lip synchronization, we leverage the widely used Sync-C and Sync-D syncnet to compute audio-visual confidence. We also incorporate hand keypoint confidence (HKC) and hand keypoint variance (HKV) lin2024cyberhost to quantify hand-pose accuracy and motion richness, respectively. For simplicity, we assume each test video contains exactly two participants: one speaking (with meaningful audio as input) and one listening (with muted audio as input). 5) Distribution distance: We employ FVD unterthiner2019fvd to measure the distance between generated and ground-truth videos.
Test Sets. We use three test sets in our experiments: 1) single-person audio conditioned human animation test set following OmniHuman lin2025omnihuman (see Tab. 1); 2) our collected two-person audio conditioned human animation test set where only one person is talking (see Tab. 1 and Tab. 4); and 3) multi-concept video customization test set following Phantom liu2025phantom , where we select 100 human-related pairs of reference images and text prompts in our experiments (see Tab. 3).
@@ 4.1 Comparisons with State-of-the-Arts
Multi-Concept Audio-Driven Human Video Generation. In Tab. 1, our approach achieves state-of-the-art or comparable performance in lip synchronization accuracy and motion diversity, particularly excelling in complex multi-person interactions where baseline methodslin2024cyberhost ; kuaishou2024 struggle with accurate audio signal assignments. For single-person scenarios, our method performs on par with specialized models like OmniHuman lin2025omnihuman . In multi-person settings, existing methods including OmniHuman and its extensions as well as leading commercial video generation models with post-processed lip-sync, fail to deliver satisfactory results. Poor lip-sync accuracy highlights their inability to generate accurate lip movements for the correct person. Although OmniHuman with oracle masks (manually setting audio-conditioned regions) improves lip-sync accuracy, it still significantly degrades overall video quality, as reflected in the FVD metric. This clearly demonstrates the limitations of existing methods that rely on single-identity assumptions, hindering their applicability to broader scenarios. In contrast, our method achieves strong performance in both lip-sync accuracy and overall video synthesis quality, effectively addressing the challenges of audio-conditioned multi-person video generation while remaining compatible with existing settings. Fig. 3 presents qualitative results for multi-person audio-driven video generation. While Kling1.6 w/ lip-sync shows many audio assignment errors and OmniHuman w/ fixed mask shows inflexible audio control with many missing cases, our method consistently assigns audio signals to the correct identity and demonstrates better motion dynamics and more precise audio-driven animation. It is worth noting that all previous methods rely on a reference frame containing complete information to generate talking-person videos, while our method only needs reference appearance of human’s head or full-body images and audios.
@@
Multi-Concept Video Customization. Although our major contribution is not on multi-concept video customization, we show that our model is also capable of preserving multi-concept visual appearances. In Tab. 3, our method outperforms existing approaches bao2024vidu ; pikaart2024 ; kuaishou2024 in preserving identity details and facial features, addressing the common degradation issue in multi-subject generation. Notably, we achieve this without sacrificing audio injection performance, while these methods are incapable to generate videos from audio conditions. It suggests our joint optimization framework on video generation and mask prediction successfully balances video generation. Our method ranks second in prompt following, likely due to our audio-driven human-centric training data focusing on talking and singing, which limits prompt diversity compared to text-to-video tasks. Despite this, qualitative results (Fig. 4) show that our approach maintains natural subject consistency and visual quality in both real-world and anime domains, outperforming previous methods that suffer from unnatural compositions and degraded visuals.
4.2 Ablation Study
We validate our local audio injection designs via ablation on three variants: 1) Global audio lin2025omnihuman applies audio across the entire feature map; 2) ID Embedding injects audio features with a learnable ID embedding without mask prediction; 3) Fixed Mask uses predefined static spatial masks. As shown in Tab. 4, our framework with predicted dynamic masks achieves the best Sync-D and FVD scores. The fixed mask yields decent Sync-D but suffers motion artifacts (worst FVD), and global audio, while scoring best in IQA, offers poor audio-visual alignment. Qualitative results in Fig. 5 further reveal that global audio drives all identities uniformly, ID embedding often mismatches audio with identities, and fixed masks lose alignment when characters move. These findings underscore that current methods overlook the need for precise local conditions, highlighting the strength of our dynamic, adaptive mask prediction strategy for multi-person talking video generation.
@@ 5 Conclusion
In this paper, we introduced InterActHuman, a novel end-to-end video diffusion framework that supports spatially aligned, multi-modal conditioning for multi-concept human animation. By integrating an efficient mask-prediction module into a pretrained DiT backbone, our method automatically infers per-identity spatio-temporal layouts from reference images and uses these masks to guide local audio conditioning. This explicit layout binding enables each reference concept to retain its unique appearance and voice, even in complex scenes with multiple interacting entities. To facilitate training procedure of our model, we presented a high-throughput pipeline for harvesting and annotating over 2.6 million identity-aware video snippets, complete with per-frame masks for humans and objects. Extensive experiments on both single- and multi-person benchmarks demonstrate that InterActHuman achieves state-of-the-art performance in lip synchronization, motion diversity and subject apperance fidelity, while maintaining competitive video quality. We hope it could serve as a solid baseline for the multi-concept human animation and audio-driven multi-person video generation community, where further improvements could be built upon it.
@@
Limitations. Since the goal of this work is for human-centric video generation, the available data domain is inherently narrower than one used for general text-to-video pretraining, limiting the ability to follow the diverse text prompts. While our framework is designed to accommodate any number of concept images, the training dataset predominantly consists of videos featuring two to three individuals. This distribution may constrain the generalization of any number of inputs, suggesting further improvements could be achieved by enlarging the diversity and scale of training data.
@@
Broader Impacts. Our model could be used to generate misinformation by using celebrity images and voices. We will strictly restrict access and add watermarks to prevent misuse.
Appendix A Algorithm of Our Model Implementation

### v2

@@ 3.3 Multi-Concept Human-Centric Video Data Curation
We curate a large-scale, human-centric video dataset from a large public video dataset (Li et al., 2024) and self-collected videos, filtering out videos that are too short (<<4 seconds) or contain too many salient humans (>>3) as detected by a human pose detector (Yang et al., 2023). Unlike previous multi-concept video generation methods (Chen et al., 2025a; Huang et al., 2025) that rely on conventional pipelines like Grounding-DINO (Liu et al., 2023) and SAM (Kirillov et al., 2023) for obtaining a single reference image per identity, our multi-stage pipeline leverages advanced vision-language models. First, each raw video is densely captioned using Qwen2-VL (Wang et al., 2024b) (distilled from Gemini-2.0-Pro), generating fine-grained descriptions of the environment, subjects’ appearance, actions, expressions, interacting salient objects, and inter-subject interactions. Next, these descriptions are parsed by a zero-shot Gemini-2.0-Flash API to extract structured appearance phrases. For spatial supervision, we use Grounding-SAM2 (Ren et al., 2024) with the query person to produce accurate, temporally consistent masks, which are used both for extracting foreground reference images (with a white background) and as ground-truth for our mask predictor. Our corpus comprises over 2.62.6M triplets of videos, per-frame masks, and captions, forming the foundation of InterActHuman.
@@ 4 Experiments
Baselines. Since ConceptMaster (Huang et al., 2025) and Video-Alchemist (Chen et al., 2025a) are not open-sourced, we do not have access to their models nor test sets. Therefore, we primarily compare our method with recent multi-concept video customization models through their publicly available APIs or public models, evaluating performance from the perspectives of visual appearance and adherence to text prompts, including Vidu2.0 (Bao et al., 2024), Pika2.1 (pik, ), Kling 1.6 (Kuaishou, 2024) and Phantom (Liu et al., 2025), and audio-driven human video generation methods, including DiffTED (Hogue et al., 2024), DiffGest (Zhu et al., 2023) + Mimiction (Zhang et al., 2024), CyberHost (Lin et al., 2024), OmniHuman (Lin et al., 2025a) and Kling 1.6 (Kuaishou, 2024) with lip synchronization.
@@
DiffTED 0.926 - 0.769 - - - -
DiffGest.+Mimic. 0.496 23.409 0.833 - - - -
CyberHost 6.627 24.733 0.884 8.974 4.011 2.856 54.797
Kling1.6 + Lip-sync. 4.449 46.490 0.826 8.401 4.716 3.444 33.555
MultiTalk - - - 7.671 4.561 3.248 35.472
OmniHuman w/o mask 7.443 47.561 0.898 9.482 4.768 3.466 33.895
OmniHuman w/ fixed mask - - - 7.068 4.690 3.369 40.239
@@
Metric Kling OmniHuman Ours Pika Phantom Kling Vidu Ours
@@
Vidu2.0 0.696 0.458 0.568 0.562 0.597 18.61 3.350 4.689
Pika2.1 0.688 0.459 0.579 0.566 0.607 19.39 3.534 4.791
Kling1.6 0.659 0.420 0.552 0.547 0.582 18.38 3.487 4.787
Phantom 0.703 0.476 0.589 0.573 0.615 17.73 3.404 4.812
@@
Evaluation Metrics. As audio and appearance are key characteristics of humans, we follow current state-of-the-art methods (Lin et al., 2025a; Huang et al., 2025) to evaluate audio-driven, multi-concept human-centric video generation. To comprehensively evaluate methods for this task, we consider five dimensions: 1) Concept fidelity: We leverage image feature extractors like CLIP-I (Radford et al., 2021) and DINO-I (Oquab et al., 2023) on subjects in the generated videos to assess whether they are aligned with the provided reference images. To get cropped subject images in output videos, we randomly sample 5 frames in each video and use Florence-2 (Xiao et al., 2024) to detect the subject bounding boxes. We then crop the detected regions and compute the CLIP-I and DINO-I scores. We also employ Face-Arc (Deng et al., 2019), Face-Cur (Huang et al., 2020) and Face-Glink (Deng et al., 2019) to evaluate the fidelity of facial features if the concept is a human. 2) Prompt following: We utilize video-level CLIP (Wang et al., 2022) to measure the similarity between the input text prompts and the resulting video content. 3) Visual quality: We utilize q-align (Wu et al., 2023a), a vision-language model, for no-reference image quality assessment (IQA) and aesthetic score estimation (AES). 4) Audio synchronization and human pose diversity: For lip synchronization, we leverage the widely used Sync-C and Sync-D (Chung & Zisserman, 2017) to compute audio-visual confidence. We also incorporate hand keypoint confidence (HKC) and hand keypoint variance (HKV) (Lin et al., 2024) to quantify hand-pose accuracy and motion richness, respectively. For simplicity, we assume each test video contains exactly two participants: one speaking (with meaningful audio as input) and one listening (with muted audio as input). 5) Distribution distance: We employ FVD (Unterthiner et al., ) to measure the distance between generated and ground-truth videos.
Test Sets. We use three test sets in our experiments: 1) single-person audio conditioned human animation test set following OmniHuman (Lin et al., 2025a) (see Tab. 1); 2) our collected two-person audio conditioned human animation test set where only one person is talking (see Tab. 1 and Tab. 4); and 3) multi-concept video customization test set following Phantom (Liu et al., 2025), where we select 100 human-related pairs of reference images and text prompts in our experiments (see Tab. 3).
@@ 4.1 Comparisons with State-of-the-Arts
Multi-Concept Audio-Driven Human Video Generation. In Tab. 1, our approach achieves state-of-the-art or comparable performance in lip synchronization accuracy and motion diversity, particularly excelling in complex multi-person interactions where baseline methods (Lin et al., 2024; Kuaishou, 2024) struggle with accurate audio signal assignments. For single-person scenarios, our method performs on par with specialized models like OmniHuman (Lin et al., 2025a). In multi-person settings, existing methods including OmniHuman and its extensions as well as leading commercial video generation models with post-processed lip-sync, fail to deliver satisfactory results. Poor lip-sync accuracy highlights their inability to generate accurate lip movements for the correct person. Although OmniHuman with oracle masks (manually setting audio-conditioned regions) improves lip-sync accuracy, it still significantly degrades overall video quality, as reflected in the FVD metric. This clearly demonstrates the limitations of existing methods that rely on single-identity assumptions, hindering their applicability to broader scenarios. In contrast, our method achieves strong performance in both lip-sync accuracy and overall video synthesis quality, effectively addressing the challenges of audio-conditioned multi-person video generation while remaining compatible with existing settings. Fig. 3 presents qualitative results for multi-person audio-driven video generation. While Kling1.6 w/ lip-sync shows many audio assignment errors and OmniHuman w/ fixed mask shows inflexible audio control with many missing cases, our method consistently assigns audio signals to the correct identity and demonstrates better motion dynamics and more precise audio-driven animation. It is worth noting that all previous methods rely on a reference frame containing complete information to generate talking-person videos, while our method only needs reference appearance of human’s head or full-body images and audios.
@@
Table 5: Runtime and parameters versus number of reference images.
Component 1 ref 2 refs 3 refs #Params
DiT Model 6.5s 7.0s 7.7s 7B
Mask Predictor 0.4s 0.8s 1.2s 56M
Full Model 6.9s 7.8s 8.9s 7B
Figure 5: Qualitative results on the failure case of mask predictions.
Multi-Concept Video Customization. Although our major contribution is not on multi-concept video customization, we show that our model is also capable of preserving multi-concept visual appearances. In Tab. 3, our method outperforms existing approaches (Bao et al., 2024; pik, ; Kuaishou, 2024) in preserving identity details and facial features, addressing the common degradation issue in multi-subject generation. Notably, we achieve this without sacrificing audio injection performance, while these methods are incapable of generating videos from audio conditions. It suggests our joint optimization framework on video generation and mask prediction successfully balances video generation. Our method ranks second in prompt following, likely due to our audio-driven human-centric training data focusing on talking and singing, which limits prompt diversity compared to text-to-video tasks. Despite this, qualitative results (Fig. 4) show that our approach maintains natural subject consistency and visual quality in both real-world and anime domains, outperforming previous methods that suffer from unnatural compositions and degraded visuals.
4.2 Ablation Study
We validate our local audio injection designs via ablation on three variants: 1) Global audio (Lin et al., 2025a) applies audio across the entire feature map; 2) ID Embedding injects audio features with a learnable ID embedding without mask prediction; 3) Fixed Mask uses predefined static spatial masks. As shown in Tab. 4, our framework with predicted dynamic masks achieves the best Sync-D and FVD scores. The fixed mask yields decent Sync-D but suffers motion artifacts (worst FVD), and global audio, while scoring best in IQA, offers poor audio-visual alignment. Qualitative results in Fig. 6 further reveal that global audio drives all identities uniformly, ID embedding often mismatches audio with identities, and fixed masks lose alignment when characters move. These findings underscore that current methods overlook the need for precise local conditions, highlighting the strength of our dynamic, adaptive mask prediction strategy for multi-person talking video generation.
Computational Cost and Runtime We measure inference time with the setting of 720p, 109 frames (28 VAE latents) on single A100, and record single forward pass time. The mask predictor adds $\sim$56M parameters versus a 7B DiT and incurs a small overhead per reference image ($\approx$0.013s per DiT block). As more references increase VAE latents from $28$ to 28+n28{+}n, DiT self-attention scales as $(1{+}n/28)^{2}$.
Failure case analysis of mask prediction In Fig. 5 we show the qualitative results on the failure case of mask predictions, where the yellow cat is the current identity image (i.e., ID query of mask predictor). We can see that although the mask prediction achieves good performance in IoU measurement from Tables 9 and 10, it still makes mistakes when there are highly overlapping regions. For example, two girls are standing in front of the yellow cat, and the mask predictor manages to exclude the upper body of the right girl in yellow because of small area of overlapping regions. Yet, the mask predictor fails for the girl in pink in both frames. In the first frame, the mask predictor manages to exclude the face region of the little girl in pink, but it mistakenly include another white animal. Besides, due to the high VAE down-sampling ratio of the widely used DiT models, the mask is predicted in a very low resolution, leading to inaccurate mask boundaries.
Additional analyses in Appendix. We include (i) a capability matrix against concurrent systems (Tables 8), (ii) mask IoU versus denoising steps/layers under low/high motion plus motion statistics (Tables 9, 10, 11), (iii) a mask cache ablation (Table 12), (iv) scaling to >>3 people (Table 13).
@@ 5 Conclusion
In this paper, we introduced InterActHuman, a novel end-to-end video diffusion framework that supports spatially aligned, multi-modal conditioning for multi-concept human animation. By integrating an efficient mask-prediction module into a pretrained DiT backbone, our method automatically infers per-identity spatio-temporal layouts from reference images and uses these masks to guide local audio conditioning. This explicit layout binding enables each reference concept to retain its unique appearance and voice, even in complex scenes with multiple interacting entities. To facilitate training procedure of our model, we presented a high-throughput pipeline for harvesting and annotating over 2.6 million identity-aware video snippets, complete with per-frame masks for humans and objects. Extensive experiments on both single- and multi-person benchmarks demonstrate that InterActHuman achieves state-of-the-art performance in lip synchronization, motion diversity and subject appearance fidelity, while maintaining competitive video quality. We hope it could serve as a solid baseline for the multi-concept human animation and audio-driven multi-person video generation community, where further improvements could be built upon it.
@@
Limitations. Since the goal of this work is for human-centric video generation, the available data domain is inherently narrower than that used for general text-to-video pretraining, limiting the ability to follow the diverse text prompts. While our framework is designed to accommodate any number of concept images, the training dataset predominantly consists of videos featuring two to three individuals. This distribution may constrain the generalization of any number of inputs, suggesting further improvements could be achieved by enlarging the diversity and scale of training data.
Acknowledgment. We thank the infrastructure team at ByteDance Intelligent Creation for the data curation and annotation support.
@@
Ethics Statement. Our model could be used to generate misinformation by using celebrity images and voices. We will strictly restrict access and add watermarks to prevent misuse.
Reproducibility Statement. We have provided a code re-implemented on the publicly available video diffusion pretraining model Wan2.1 (Wang et al., 2025a) to show the details of our method. We also provided the dataset processing code, e.g., caption analysis and Grounding-SAM2 (Ren et al., 2024). In the Algorithm 1, we provided detailed pseudo-code to show the inference pipeline of our method. Based on such information, we believe the reproducibility of method is solid enough.
Appendix A Additional Experimental Results
A.1 LLM Usage Statement.
We use LLMs (e.g., Gemini-2.5 and GPT-5) to polish our paragraphs.
A.2 Comparison with Single-Person Talking-Head Methods
Although our method targets multi-person full-body talking video generation, we report single-person talking-head results for completeness.
Table 6: CelebV-HQ: higher is better for IQA/ASE/Sync-C; lower is better for FID.
[…]
SadTalker (Zhang et al., 2023) 2.953 1.812 3.843 36.648
Hallo (Xu et al., 2024a) 3.505 2.262 4.130 35.961
VExpress (Wang et al., 2024a) 2.946 1.901 3.547 65.098
EchoMimic (Chen et al., 2024) 3.307 2.128 3.136 35.373
Hallo-3 (Cui et al., 2024) 3.451 2.257 3.933 38.481
Loopy (Jiang et al., 2024a) 3.780 2.492 4.849 33.204
OmniHuman (Lin et al., 2025a) 3.875 2.656 5.199 31.435
Ours 3.834 2.553 5.047 30.159
As shown in Table 6, OmniHuman achieves the best single-person talking-head scores on IQA/ASE/Sync-C, while our method yields the best FID. Since our approach targets multi-person, full-body generation rather than single-person heads, these results indicate competitive quality under a setting that is not our primary target. On RAVDESS (Table 7), our method attains the best IQA and ASE, while OmniHuman leads on Sync-C and FID. This mirrors the CelebV-HQ trend and supports our claim that the proposed framework remains competitive on single-person benchmarks while being designed for multi-person, full-body scenarios.
Table 7: RAVDESS: higher is better for IQA/ASE/Sync-C; lower is better for FID.
[…]
SadTalker (Zhang et al., 2023) 3.840 2.277 4.304 32.343
Hallo (Xu et al., 2024a) 4.393 2.688 4.062 19.826
VExpress (Wang et al., 2024a) 3.690 2.331 5.001 26.736
EchoMimic (Chen et al., 2024) 4.504 2.742 3.292 21.058
Hallo-3 (Cui et al., 2024) 4.006 2.462 4.448 28.840
Loopy (Jiang et al., 2024a) 4.506 2.658 4.814 17.017
OmniHuman (Lin et al., 2025a) 4.564 2.815 5.255 16.970
Ours 4.602 2.915 5.132 17.187
A.3 Comparison with concurrent works
We show comparison with recent multi-person talking systems which are conditioned on a reference frame only and released concurrently with our method. Our framework supports both concept-only and first-frame modes.
Table 8: Qualitative capability comparison. $\checkmark$: supported; x: not supported.
Methods Audio-driven full-body First-frame I2V Multi-ref images Multi-person talking
Video-Alchemist (Chen et al., 2025a) x x $\checkmark$ x
ConceptMaster (Huang et al., 2025) x x $\checkmark$ x
Phantom (Liu et al., 2025) x x $\checkmark$ x
OmniHuman (Lin et al., 2025a) $\checkmark$ $\checkmark$ x x
HunyuanVideo-Avatar (Chen et al., 2025b) $\checkmark$ $\checkmark$ x $\checkmark$
MultiTalk (Kong et al., 2025) $\checkmark$ $\checkmark$ x $\checkmark$
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
A.4 Mask Refinement Across Denoising Steps and Layers
Across both regimes (Tables 9, 10), IoU steadily improves with denoising steps, validating the iterative refinement strategy. Deeper layers (e.g., layer 36) outperform shallow layers early on, while the combined mask yields the strongest mid/late-step IoUs. Notably, high-motion sequences retain strong IoUs and achieve competitive lip-sync (Sync-D 6.921), indicating that masks are not static and remain reliable even under larger motion (see the motion statistics in Table 11).
Table 9: Low motion strength: mask IoU across steps/layers; Sync-D for the setting.
[…]
Mask layer 4 0.306 0.436 0.509 0.640 0.858
Mask layer 20 0.386 0.629 0.733 0.890 0.957
Mask layer 36 0.538 0.742 0.850 0.916 0.923
Combine mask 0.376 0.738 0.881 0.949 0.956 7.292
Table 10: High motion strength: mask IoU across steps/layers; Sync-D for the setting.
[…]
Mask layer 4 0.113 0.426 0.529 0.730 0.779
Mask layer 20 0.527 0.816 0.911 0.934 0.945
Mask layer 36 0.694 0.902 0.923 0.931 0.915
Combine mask 0.741 0.916 0.932 0.936 0.937 6.921
Table 11: Statistics used to define motion regimes.
Statistic Low High
Body keypoints confidence 0.737 0.663
Max body kp moving dist (px) 21.268 216.269
Var body kp moving dist (px) 6.098 34.022
Head keypoints confidence 0.9856 0.9840
Max head kp moving dist (px) 2.258 34.201
Var head kp moving dist (px) 2.961 216.383
A.5 Effect of Mask Cache
Using the cached mask from step t1t{-}1 to gate audio injection at step $t$ markedly improves lip sync (Table 12; 6.921 vs 11.046 Sync-D). Without caching, identity assignment degrades toward a single-person behavior, confirming the necessity of cross-timestep spatial guidance.
Table 12: Mask cache significantly improves multi-person lip sync.
[…]
With mask cache 6.921
Without mask cache 11.046
A.6 Scaling Beyond Three People
Table 13 shows stable scaling to 4–5 speakers: Sync-D improves slightly (6.608) and AES increases (3.992) with a marginal IQA change. This aligns with our claim that per-identity mask prediction is independent, enabling more entities without collapsing lip-sync quality.
Table 13: Stable performance when scaling to 4–5 subjects.
[…]
$\leq$3 6.670 3.467 4.757
4–5 6.608 3.992 4.738
Appendix B Algorithm of Our Model Implementation